OpenStudy (fantalux):

Need Help with turning equations into standard form! Can anyone help, will give medals!!!

OpenStudy (fantalux):

Which shows this equation in standard form? 6 – 2(x – y) = 3x + 5 A. 5x – 2y = 1 B. 5x – 2y = –1 C. 5x + 2y = 1 D. 2x – 5y = 1

OpenStudy (fantalux):

I know i have to subtract y from both sides but where do i plug it in?

OpenStudy (bibby):

Start by distributing the 2. 6 – 2(x – y) = 3x + 5 6-[2x-y] 6-2x+y=3x+5

OpenStudy (fantalux):

Thankyou Bibby!!!!

OpenStudy (bibby):

We're not done yet Fanta!!!

OpenStudy (fantalux):

Oh lol....

OpenStudy (bibby):

Anyhow.... :D 6-2x+2y=3x+5 *** I messed up the subtraction

OpenStudy (fantalux):

...I got 4x+2y=8x......

OpenStudy (bibby):

6-2x+2y=3x+5 *** add 2x to both sides 6+2y=5x+5 subtract 5 from both sides 1+2y=5x

OpenStudy (fantalux):

So the answer is 5x+2y=1???

OpenStudy (bibby):

No. try manipulating\[\large 1+2y=5x\] into one of the equations

OpenStudy (fantalux):

ummmm....... 5x-2y=1 or does the 2 have to me negative???

OpenStudy (bibby):

That's right. if you subtract 2y from both sides you get 5x-2y = 1 which is one of the answers

OpenStudy (fantalux):

Ya! Lol Thankyou Bibby!